OpenStudy (anonymous):

Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT.

OpenStudy (anonymous):

@iPwnBunnies im back lol

OpenStudy (ipwnbunnies):

Ok.

OpenStudy (somy):

<.< and values?

OpenStudy (ipwnbunnies):

Somy c:

OpenStudy (anonymous):

sorry had to grab lunch im typing them now

OpenStudy (somy):

Bunnies♥

OpenStudy (kainui):

kanuey <3

OpenStudy (ipwnbunnies):

Kainui c:

OpenStudy (anonymous):

m = 27.776 c = i dont exactly know this one because i was not given the heat capacity. delta T = 100.5 - 38.9

OpenStudy (somy):

Kai ♥♥♥

OpenStudy (anonymous):

and i assume q water is what we are finding

OpenStudy (somy):

dude @Avenged7x you have all the values and formula >.>

OpenStudy (somy):

Specific heat capacity of water is 4.18

OpenStudy (somy):

its i guess constant value for water :D

OpenStudy (kainui):

I thought heat capacity of water was defined to be 1 or something. I don't know, maybe I'm in the wrong units, it's been way too long.

OpenStudy (anonymous):

i did this similar on a test and failed it 10 questions just like these thats why i am getting this done so somebody can help me find what i did wrong. while also getting another assignment done :)

OpenStudy (somy):

you know the change in temperature u know mass and C= 4.18

OpenStudy (ipwnbunnies):

That's density, Kai. 1 kg/m^3

OpenStudy (somy):

u took C as what? @Avenged7x

OpenStudy (somy):

yeah lol @Kainui sleepy head ♥

OpenStudy (anonymous):

well when you said 4.18 i remeber bunnies told me to always remember heat capacity of water is 4.179/g celsius. but that was one problem i couldnt figure out what c was until just now. when you reminded me

OpenStudy (kainui):

Actually I am right. \[\Large C_{water}=1\frac{ calorie }{gram \ degree C}\]

OpenStudy (kainui):

Water is so common it's the standard for a lot of stuff lol =P

OpenStudy (anonymous):

so that means it will look like this: 27.776 x 4.179 x 100.5 - 38.9?

OpenStudy (somy):

yeah basically

OpenStudy (anonymous):

is there anything else i have to do other than solving it out right like this?

OpenStudy (anonymous):

*like it reads

OpenStudy (somy):

@Kainui dk this one

OpenStudy (somy):

nope, its only this much @Avenged7x

OpenStudy (somy):

what was unit of you mass tho

OpenStudy (somy):

coz it depend on it if mass is in Kg, c= 4181 if mass is in g, c=4.18

OpenStudy (anonymous):

it is grams

OpenStudy (somy):

right good then :D

OpenStudy (anonymous):

ok doing the problem as it reads i got 8093.04... is that right?

OpenStudy (somy):

wrong

OpenStudy (somy):

dear its like this Q= mc Delta T = 27.776 * 4.179 * (100.5-38.9)

OpenStudy (anonymous):

thats numerical part is what i put into my calculator. and i got 8093.04

OpenStudy (anonymous):

nm i got 7150.28

OpenStudy (anonymous):

i have a ti nspire cx cas so the parenthesis made a difference

OpenStudy (somy):

lol yeah second one is right :)

OpenStudy (anonymous):

ok thanks that makes sense now heres a medal!

OpenStudy (somy):

:)